# OUJ-FTC-15

### v4

@@ 3 The model of the SM with toponium
In Figure 1, for the case with $M_{t}=172.69\ \mathrm{GeV},\ \alpha_{s}(M_{Z})=0.1189$, and $\Lambda=400\ \mathrm{GeV}$, the behavior is plotted with the energy scale on the $x$-axis and the effective potential on the $y$-axis. It can be seen that the minimum value of the term for $\lambda(\mu)$ including quantum corrections goes below about 0.15-0.15. This indicates that the vacuum stability is deteriorating. This deterioration is assessed relative to the SM, where $\lambda(\mu)$ with quantum corrections included in the RGEs reaches approximately 0.025-0.025 at one loop and 0.015-0.015 at three loops[16, 17]. Such a substantially more negative behavior is incompatible with the observed longevity of the Universe (i.e., a long-lived Electroweak vacuum) and is therefore problematic[20, 21].
@@ 5 Summary and Discussion
These results indicate that the vacuum stability can be improved when an additional field is incorporated into the SM. However, when toponium is introduced into the SM, the vacuum stability is instead worsened in both the one-loop effective potential and RGEs. This is in comparison with the case in the SM, where $\lambda(\mu)$ with quantum corrections included in the RGEs reaches approximately 0.025-0.025 at one loop and about 0.015-0.015 at three loops[16, 17]. This result is incompatible with the observational fact that the Universe has persisted to the present day[20, 21]. Further investigations including higher-loop corrections are essential, and it is also necessary to experimentally evaluate the validity of the observed excess in $t\bar{t}$ events. As an immediate extension, it will be important to carry out a similar study within the framework of the Two-Higgs-Doublet Models (2HDM)[22, 23, 24]. In any case, the results reported by the LHC are highly intriguing and may open a path toward New Physics.

### v5

@@ 3 The model of the SM with toponium
In Figure 1, for the case with $M_{t}=172.69\ \mathrm{GeV},\ \alpha_{s}(M_{Z})=0.1189$, and $\Lambda=400\ \mathrm{GeV}$, the behavior is plotted with the energy scale on the $x$-axis and the effective potential on the $y$-axis. It can be seen that the minimum value of the term for $\lambda(\mu)$ including quantum corrections goes below about 0.04-0.04. This indicates that the vacuum stability is deteriorating. This deterioration is assessed relative to the SM, where $\lambda(\mu)$ with quantum corrections included in the RGEs reaches approximately 0.025-0.025 at one loop and 0.015-0.015 at three loops[16, 17]. Such a substantially more negative behavior may be incompatible with the observed longevity of the Universe (i.e., a long-lived Electroweak vacuum) and is therefore problematic[20, 21].
@@ 5 Summary and Discussion
These results indicate that the vacuum stability can be improved when an additional field is incorporated into the SM. However, when toponium is introduced into the SM, the vacuum stability is instead worsened in both the one-loop effective potential and RGEs. This is in comparison with the case in the SM, where $\lambda(\mu)$ with quantum corrections included in the RGEs reaches approximately 0.025-0.025 at one loop and about 0.015-0.015 at three loops[16, 17]. This result may be incompatible with the observational fact that the Universe has persisted to the present day[20, 21]. Further investigations including higher-loop corrections are essential, and it is also necessary to experimentally evaluate the validity of the observed excess in $t\bar{t}$ events. As an immediate extension, it will be important to carry out a similar study within the framework of the Two-Higgs-Doublet Models (2HDM)[22, 23, 24]. In any case, the results reported by the LHC are highly intriguing and may open a path toward New Physics.